OpenStudy (anonymous):

A kite 50ft above the ground moves horizontally at a speed of 8ft/s. At what rate is the angle between the string and the horizontal decreasing when 300ft of string has been let out?

OpenStudy (anonymous):



OpenStudy (anonymous):

tan(theta) = 50/x differentiate both sides with respect to time find x when sqrt(x^2+50^2) =300 (*spoiler*) find theta when x= sqrt(300^2 -50^2) plug x, theta and dx/dt into the result of the differentiation to find d(theta)/dt

OpenStudy (anonymous):

can you explain the diferrentiatie both sides with respect to time part?

OpenStudy (anonymous):

@Algebraic!

OpenStudy (anonymous):

I guess, it's just implicit differentiation really...or the chain rule... or whatever you are comfortable thinking of it as... d/dt ( tan ( f(t) ) =

OpenStudy (anonymous):

derivative of the 'outside' is...?

OpenStudy (anonymous):

sec^2

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so it's x = 50/sec^2(theta)?

OpenStudy (anonymous):

yep, so sec^2 (f(t) ) * f '(t)

OpenStudy (anonymous):

f(t) is theta yeah, I wrote it that way because theta depends on time and I wanted you to see the chain rule

OpenStudy (anonymous):

so LHS is sec^2(theta) * d(theta) /dt RHS still needs to be differentiated...

OpenStudy (anonymous):

d/dt ( 1/(g(t)) ) =...?

OpenStudy (anonymous):

-1/ ..........

OpenStudy (anonymous):

what is g(t) in this case?

OpenStudy (anonymous):

well, just because I used f(t) for theta... now we're talking about x, which is a different function of time...

OpenStudy (anonymous):

ah okay

OpenStudy (anonymous):

-1/x^2*dx/dt

OpenStudy (anonymous):

great:)

OpenStudy (anonymous):

so:\[\frac{ d \theta }{dt } = \frac{ -1 }{x^2 \sec^2\theta }\frac{ d x}{ dt }\]

OpenStudy (anonymous):

dx/dt is given.. x and theta are easy to find...

OpenStudy (anonymous):

dx/dt = 8

OpenStudy (anonymous):

d(theta)/dt = -8/x^2*sec(theta)

OpenStudy (anonymous):

x= sqrt(300^2 -50^2) theta = arctan (50/x)

OpenStudy (anonymous):

wouldn't x = sqrt(300^2 + 50^2)?

OpenStudy (anonymous):

x = 304.138

OpenStudy (anonymous):

theta = .1629 rads

OpenStudy (anonymous):

@Algebraic!

OpenStudy (anonymous):

so x is longer than the hypotenuse?

OpenStudy (anonymous):

x = 295.803

OpenStudy (anonymous):

k:)

OpenStudy (anonymous):

so d(theta)/dt = -8/((295.803)^2*(sec^2(.16744 rads)) which is incorrect.

OpenStudy (anonymous):

-8.87E-5?

OpenStudy (anonymous):

well you said d(theta)/dt = -1/(x^2*sec^2(theta)) * dx/dt

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you didn't get -8.87E-5?

OpenStudy (anonymous):

yes I did

OpenStudy (anonymous):

and it says it's wrong...

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

dunno let me look it all over. I don't see any glaring mistakes...

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

-8.89E-5 rad.s/sec .. best I can do... let's see who else is on, might be able to see if I made a mistake...

OpenStudy (anonymous):

May I have the answer to the question?

OpenStudy (anonymous):

go for it.

OpenStudy (anonymous):

@eSpeX

OpenStudy (anonymous):

@RolyPoly id like the answer too!

OpenStudy (anonymous):

@callisto is checking it over...

OpenStudy (anonymous):

ty @Callisto

OpenStudy (callisto):

I... am no good at maths... \[tan \theta = \frac{50}{x}\] Differentiate both sides with respect to x. Probably you won't get: \[\frac{ d \theta }{dt } = \frac{ -1 }{x^2 \sec^2\theta }\frac{ d x}{ dt }\]

OpenStudy (callisto):

My bad, I meant with respect to t

OpenStudy (anonymous):

you're saying the differentiation is wrong?

OpenStudy (callisto):

\[tan \theta = \frac{50}{x}\] Diff. both sides w.r.t. t \[sec^2\theta \frac{d\theta}{dt} = \frac{-50}{x^2} \frac{dx}{dt}\]\[\frac{d\theta}{dt} = \frac{-50}{x^2sec^2\theta } \frac{dx}{dt}\]

OpenStudy (anonymous):

arg

OpenStudy (anonymous):

forgot the 50 rfl

OpenStudy (anonymous):

you so smart cally

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

-50*-8/((295.803^2)*sec^2(.16744)) = correct answer

OpenStudy (anonymous):

thank you @Callisto and @Algebraic!

OpenStudy (anonymous):

yes thanks @Callisto !

OpenStudy (callisto):

\[sec^2\theta=\frac{x^2+50}{x^2}\]You can cancel the x^2 since \[\frac{d\theta}{dt} = \frac{-50}{x^2sec^2\theta } \frac{dx}{dt}=\frac{-50}{x^2(\frac{x^2+50}{x^2}) } \frac{dx}{dt}\]\[=\frac{-50}{(x^2+50) } \frac{dx}{dt}\]And x^2 is easy to find

OpenStudy (anonymous):

only took me 4hrs but i got my question worth 1pt right now to get some sleep

OpenStudy (anonymous):

heh

OpenStudy (anonymous):

you guys have a good night thanks again

OpenStudy (callisto):

Good night!!~

OpenStudy (anonymous):

A revised Mathematica solution, posted on 25 October 2012, 22:15 California time, is attached.